# Stable Signer: Hierarchical Sign Language Generative Model

### v1

@@ 1 Introduction
In the field of sign language based on deep learning, Sign Language Recognition (SLR hu2023continuous; skeleton_aware_slr; slt-how2sign-wicv2023) first developed and became popular, and then around 2015-2020, Sign Language Production (SLP SLP_Mixed; SLP_ProgTransf; Fang_2025_ICCV) also gradually became popular. However, compared to the mature and popular SLR and understanding tasks, sign language production is relatively lagging behind. This is caused by several reasons: (1) The processing methods for different datasets vary, and there are differences in accuracy, which makes the learning of postures relatively difficult. (2) Different generation or evaluation models are specific to certain posture formats, making the comparison and transfer relatively difficult. (3) The multi-stage SLP pipeline suffers from error accumulation across stages, making performance improvements challenging.
@@
To address these issues, drawing on some of the similar work shazeer2017outrageouslylargeneuralnetworks; riquelme2021scalingvisionsparsemixture in Sign Language Production (SLP), we realized that there is redundancy in the process of SLP based on deep learning: the input and output of the SLP task are text and the final high-quality video. For pose video, rules can be used as a prior Stoll2020text2sign, because the gloss and the text only need to correspond to a specific gesture or movement. The SLR cannot introduce rule priors because there are multiple possible inputs of similar gestures, while in SLP, as long as the precise gesture output is required, the core problem of SLP is actually that the input of the gloss needs to be accurate, so the pressure lies in the language understanding NEURIPS2022_ec795aea part.
And what’s more serious is that the SLP methods actually did not take into account that their requirements are different from those of SLR and SLT. As shown in Fig 1, for a specific posture, there are multiple expressions. If we forcibly learn the generation of the posture, we are likely to fall into the average posture (suboptimal state). If our training goals or data are not always precisely aligned, the more postures there are, the less accurate or difficult it is to learn at same time. Therefore, the motivation of end-to-end and the two existing problems prompt us to streamline this redundant part saunders2020progressivetransformersendtoendsign; saunders2021continuous3dmultichannelsign. Then we can focus on researching and improving the truly necessary parts of language understanding NEURIPS2022_ec795aea and Pose2Vid.
Since we can now learn SLP in an rarely end-to-end manner, we need to consider the challenges that text understanding and pose-conditioned video generation will face: (1) Traditional SLP uses Gloss or Text as input inproceedings. To ensure accuracy, they also use translation models to translate Prompts or complex Text into Gloss, but it is not as thorough as ours. (2) When dealing with processing, many previous methods, in order to handle heterogeneous sign language data, repeatedly train the model or specially create methods, which unnecessarily increases the complexity of the task. (3) For the Pose2Video chan2019everybodydance; saunders2020everybodysignnowtranslating part, previous methods, due to the lack of end-to-end training, have difficulty using complex textual information for assistance or combining the pose input of Pose2Vid model more closely. This is the main reason why it is difficult to improve the final video effect.
@@
Figure 2: Overview of the Prompt2Pose and Pose2Video pipeline: (a) Text/Prompt to Gloss: The SLUL module uses a T5 encoder to process complex prompts and generate precise gloss sequences, trained with SLUL Loss, SAGM Loss, KL divergence, and contrastive loss. (b) Gloss to Pose: A rule-based stable pose prior database provides candidate poses, which are selected by the Pose Chosen Networks (MLP) guided by semantic features. The Gating Network and Expert Networks in SLP-MoE determine the optimal pose selection with stability constraints applied. (c) Pose to Video: The stabilized pose sequence is fed into a diffusion renderer to generate high-quality sign language videos in multiple styles, with the semantic features reused for fine-tuning the MoE module.
@@ 2.1 Data Construction
For the sign language, we use a dataset called Prompt2Sign Fang_2025_ICCV. The videos in this dataset are sourced from existing mainstream datasets and online sign language videos, and they are segmented into tens of thousands of clips. It has LLM-generated prompts, texts, and comprehensive glosses. It also provides a unified processing procedure to facilitate our comparison with existing methods. For pose condition control, we first use a video dataset called OpenVidHD [^3] https://github.com/NJU-PCALab/OpenVid-1M nan2025openvid1mlargescalehighqualitydataset for training to improve video quality. Then, we use sign language videos to enhance the posture guidance ability.
@@ 2.3 SLP-MoE for Stable Pose2Video Production
NSLP-G hwang2021non - - - - - 5.75 8.21 11.62 17.55 31.98
Fast-SLP Transformers 11099440 17.19 23.11 29.49 36.96 55.85 12.85 17.35 23.38 39.46 46.89
Neural Sign Actors Baltatzis_2024_CVPR - - - - - 13.12 18.25 25.44 41.31 47.55
SignLLM-1x1B-Super-P (ASL) Fang_2025_ICCV 18.68 25.11 31.99 40.14 60.47 13.93 18.86 25.40 42.87 50.91
@@
Table 2: Ablation Study: A comparison of the generation performance under different settings of our model. The first three lines represent the generation of Pose Video, and the last line shows the final video generated by SLP MoE. We compared the impact of changes in different stages on the entire process, and found that our changes significantly improved the overall generation effect. Base: We use the base model designed by T5 2020t5. SLUL: Sign Language Understanding Linker. SLP-MoE: Sign Language Production Mixture-of-Experts. SAGM Loss: Semantic-Aware Gloss Masking Loss. The results show that all the changes were beneficial.
@@ 3.1 Experimental Setup
As mentioned above, our training is based on the Prompt2Sign dataset Fang_2025_ICCV and the American Sign Language (ASL) portion of WLASL WLASL. They have 30k and 10k video clips of ASL respectively, and some of the missing parts (Prompts and Glosses) were properly completed or deleted.
@@ 3.2 Evaluation for Sign Language Production
In Table 3, we present final video generation results with SLP-MoE. Compared to Fast-SLP Transformers (SignDiffusion 11099440), we achieve 48.6% improvement on BLEU-4 (dev) and 63.7% on test set. Our test BLEU-4 (21.03) surpasses latest Neural Sign Actors (13.12) Baltatzis_2024_CVPR by a large margin, validating our hierarchical generation approach.
@@
Neural Sign Actors Baltatzis_2024_CVPR 13.12 47.55
Fast-SLP Transformers 11099440 17.19 55.85 12.85 46.89
@@
Ablation Study. We conduct ablation studies on the How2Sign dataset to validate each component’s contribution, as shown in Table 2. We evaluate progressive improvements from base T5 2020t5 through SLUL, SAGM Loss, and SLP-MoE additions on both development and test sets.
@@
Zhang 10.1145/3706598.3713855 - - 27.60 66.40
SignLLM Fang_2025_ICCV 23.10 58.76 22.05 56.46
@@
We compare SLUL’s Prompt2Gloss performance with previous Text2Gloss approaches in Table 4. This is important as users naturally input prompts like “How do you sign ‘I carried it’?” rather than simplified text. Our approach achieves test BLEU-4 of 30.74 and ROUGE of 69.72, representing 43.3% and 23.5% improvements over SignLLM Fang_2025_ICCV. Despite handling more complex prompts than baseline Text2Gloss tasks, we outperform all methods including 10.1145/3706598.3713855 by 11.4% on test BLEU-4. This demonstrates that SAGM Loss effectively handles prompt nuances including ambiguity and implicit semantics. Consistent performance across dev and test sets indicates robust generalization rather than overfitting.
@@ 3.3 Evaluation for Video Generation Quality
The simplified pose targets shown in Figure 5(a) demonstrate a critical innovation of our approach—the model automatically learns to generate clean, semantically meaningful pose representations through the synergistic integration of rule-based pose priors and SLUL-based semantic guidance. These simplified poses effectively eliminate the noise, jitter, and temporal instability commonly present in raw pose estimations from conventional video analysis tools such as DWPose yang2023effective. Of course, what is even more important is as shown in the first row of the picture: different sign language users have different expressions. These gestures are not aligned in time and space, which causes the models that previously learned average gestures to often be unstable, or to generate restricted or unconfident gestures. However, our method will automatically learn deterministic gestures. As long as one of the multiple gestures is completed, it is considered correct. Therefore, our generation part is more stable.
@@
Stoll et al. stoll2020signsynth 0.727 0.533 23.17 64.01
SignGAN Saunders_2022_CVPR 0.759 0.605 22.05 27.75
SignDiff 11099440 0.849 0.676 20.04 25.22
@@
Zhang 10.1145/3706598.3713855 0.864 - - 56.28
SignDiff 11099440 0.849 0.676 20.04 25.22
@@
Figure 6: Ablation Study on SLP-MoE’s Training Efficiency: We examined the efficiency under the condition where there was no SLP-MoE (using the regular Pose2Vid model instead cheng2025wan). The results showed that using our SLP-MoE for joint training could achieve more thorough convergence, train faster, and produce better results.
@@ A.1 Sign Language Production
Sign Language Production (SLP) aims to generate sign language videos or motion sequences from spoken language inputs, bridging communication gaps for the deaf and hard-of-hearing community gregory2013bridgingcommunicationgapswiththedeaf. Early deep learning approaches decomposed SLP into cascaded steps: Text-to-Gloss (T2G) abdullah2025stateofthearttranslationtexttoglossusing, Gloss-to-Pose (G2P) xie2023g2pddmgeneratingsignpose, and Pose-to-Sign (P2S) moryossef2025posebasedsignlanguageappearance. However, these pipeline methods struggled with smooth motion blending between signs. Recent works have explored end-to-end approaches using transformer architectures and diffusion models. Neural Sign Actors baltatzis2024neuralsignactorsdiffusion proposed a diffusion-based model trained on 4D signing avatars with anatomically informed graph neural networks defined on the SMPL-X skeleton for realistic 3D sign production. T2S-GPT yin2024t2sgptdynamicvectorquantization introduced dynamic vector quantization for autoregressive SLP, while SignLLM Fang_2025_ICCV extended SLP to a multilingual framework supporting eight sign languages through novel reinforcement learning components. Despite these advances, generating realistic, semantically accurate signing motions with fine-grained hand and facial expressions remains challenging walsh2025slrtp2025signlanguageproduction; TII-Web-2023-Falcon; Cho-arxiv-2022-Prompt; fang2025signxfoundationmodelsign; hwang2021non.
@@ A.2 Motion Generation
Human motion generation has witnessed significant progress with the adoption of diffusion models. MotionDiffuse zhang2022motiondiffusetextdrivenhumanmotion was among the first diffusion-based frameworks for text-driven motion generation, demonstrating probabilistic mapping through denoising steps and enabling fine-grained body part control. MDM (Motion Diffusion Model) tevet2022humanmotiondiffusionmodel introduced a transformer-based architecture with classifier-free guidance, predicting motion samples rather than noise to facilitate geometric losses on joint positions and velocities. LGTM Sun_2024 proposed a Local-to-Global pipeline leveraging large language models to decompose motion descriptions into part-specific narratives, addressing local semantic accuracy issues. Recent works have also explored music-to-dance synthesis, compositional action generation liu2024languagefreecompositionalactiongeneration, and stylized motion control zhong2025smoogptstylizedmotiongeneration. These advances provide foundational techniques for generating temporally coherent human movements, which are directly applicable to sign language motion synthesis where precise hand articulation and body coordination are essential.
Figure 7: The framework of Wan-Animate cheng2025wan. It utilizes a unified input paradigm and a decoupled control strategy. Spatially-aligned skeleton signals are directly injected into the DiT backbone for body control, while facial expressions are driven by implicit features extracted from reference images and injected via cross-attention layers.
@@ A.3 Controllable Video Generation
Controllable video generation has emerged as an active research direction, extending image-based control mechanisms to the temporal domain. ControlNet zhang2023addingconditionalcontroltexttoimage pioneered the integration of structural controls such as depth maps, edge detection, and human pose into text-to-image diffusion models through zero convolution layers. ControlVideo zhang2023controlvideotrainingfreecontrollabletexttovideo adapted ControlNet to video generation without fine-tuning, inheriting high-quality consistent frame generation through fully cross-frame interaction and interleaved-frame smoothing. ControlNeXt peng2025controlnextpowerfulefficientcontrol further improved efficiency by reducing trainable parameters up to 90% while supporting diverse conditional controls including pose sequences for video diffusion. Follow-Your-Pose ma2024followposeposeguidedtexttovideo and MimicMotion zhang2025mimicmotionhighqualityhumanmotion specifically addressed human video generation with pose guidance. Most recently, Wan-Animate cheng2025wan advances this field by leveraging a Diffusion Transformer (DiT) peebles2023scalablediffusionmodelstransformers; tevet2022humanmotiondiffusionmodel; fang2025streamflowtheoryalgorithmimplementation; NEURIPS2022_ec795aea architecture. As illustrated in Figure 7, it employs a decoupled control paradigm that integrates spatially-aligned skeleton signals for body movements and implicit facial features via cross-attention for fine-grained expression reenactment. These controllable generation techniques offer promising directions for sign language video synthesis SLcomprehensibility; signor; Xu-SIGPLAN-2022-Systematic, where pose-conditioned generation can ensure anatomically correct and semantically aligned signing motions.

### v2

@@ 1 Introduction
In the field of sign language based on deep learning, Sign Language Recognition (SLR Hu et al. (2023); skeleton_aware_slr; Tarrés et al. (2023)) first developed and became popular, and then around 2015-2020, Sign Language Production (SLP SLP_Mixed; SLP_ProgTransf; Fang et al. (2025a)) also gradually became popular. However, compared to the mature and popular SLR and understanding tasks, sign language production is relatively lagging behind. This is caused by several reasons: (1) The processing methods for different datasets vary, and there are differences in accuracy, which makes the learning of postures relatively difficult. (2) Different generation or evaluation models are specific to certain posture formats, making the comparison and transfer relatively difficult. (3) The multi-stage SLP pipeline suffers from error accumulation across stages, making performance improvements challenging.
@@
To address these issues, drawing on some of the similar work Shazeer et al. (2017); Riquelme et al. (2021) in Sign Language Production (SLP), we realized that there is redundancy in the process of SLP based on deep learning: the input and output of the SLP task are text and the final high-quality video. For pose video, rules can be used as a prior Stoll2020text2sign, because the gloss and the text only need to correspond to a specific gesture or movement. The SLR cannot introduce rule priors because there are multiple possible inputs of similar gestures, while in SLP, as long as the precise gesture output is required, the core problem of SLP is actually that the input of the gloss needs to be accurate, so the pressure lies in the language understanding Saharia et al. (2022) part.
And what’s more serious is that the SLP methods actually did not take into account that their requirements are different from those of SLR and SLT. As shown in Fig 1, for a specific posture, there are multiple expressions. If we forcibly learn the generation of the posture, we are likely to fall into the average posture (suboptimal state). If our training goals or data are not always precisely aligned, the more postures there are, the less accurate or difficult it is to learn at same time. Therefore, the motivation of end-to-end and the two existing problems prompt us to streamline this redundant part saunders2020progressivetransformersendtoendsign; saunders2021continuous3dmultichannelsign. Then we can focus on researching and improving the truly necessary parts of language understanding Saharia et al. (2022) and Pose2Vid.
Since we can now learn SLP in an rarely end-to-end manner, we need to consider the challenges that text understanding and pose-conditioned video generation will face: (1) Traditional SLP uses Gloss or Text as input Zelinka and Kanis (2020). To ensure accuracy, they also use translation models to translate Prompts or complex Text into Gloss, but it is not as thorough as ours. (2) When dealing with processing, many previous methods, in order to handle heterogeneous sign language data, repeatedly train the model or specially create methods, which unnecessarily increases the complexity of the task. (3) For the Pose2Video Chan et al. (2019); Saunders et al. (2020) part, previous methods, due to the lack of end-to-end training, have difficulty using complex textual information for assistance or combining the pose input of Pose2Vid model more closely. This is the main reason why it is difficult to improve the final video effect.
@@
Figure 2: Overview of the Prompt2Pose and Pose2Video pipeline: (a) Text/Prompt to Gloss: The SLUL module uses a T5 encoder to process complex prompts and generate precise gloss sequences, trained with SLUL Loss, SAGM Loss, KL divergence, and contrastive loss. (b) Gloss to Pose: A rule-based stable pose prior database provides candidate poses, which are selected by the Pose Chosen Networks (MLP) guided by semantic features and trained to ensure natural and smooth pose transitions. (c) Pose to Video: The stabilized pose sequence is fed into a diffusion renderer to generate high-quality sign language videos in multiple styles, with the semantic features reused for fine-tuning the MoE module.
@@ 2.1 Data Construction
For the sign language, we use a dataset called Prompt2Sign Fang et al. (2025a). The videos in this dataset are sourced from existing mainstream datasets and online sign language videos, and they are segmented into tens of thousands of clips. It has LLM-generated prompts, texts, and comprehensive glosses. It also provides a unified processing procedure to facilitate our comparison with existing methods. For pose condition control, we first use a video dataset called OpenVidHD [^3] https://github.com/NJU-PCALab/OpenVid-1M Nan et al. (2025) for training to improve video quality. Then, we use sign language videos to enhance the posture guidance ability.
@@ 2.3 SLP-MoE for Stable Pose2Video Production
NSLP-G Hwang et al. (2021) - - - - - 5.75 8.21 11.62 17.55 31.98
Fast-SLP Transformers Fang et al. (2025c) 17.19 23.11 29.49 36.96 55.85 12.85 17.35 23.38 39.46 46.89
Neural Sign Actors Baltatzis et al. (2024a) - - - - - 13.12 18.25 25.44 41.31 47.55
SignLLM-1x1B-Super-P (ASL) Fang et al. (2025a) 18.68 25.11 31.99 40.14 60.47 13.93 18.86 25.40 42.87 50.91
@@
Table 2: Ablation Study: A comparison of the generation performance under different settings of our model. The first three lines represent the generation of Pose Video, and the last line shows the final video generated by SLP MoE. We compared the impact of changes in different stages on the entire process, and found that our changes significantly improved the overall generation effect. Base: We use the base model designed by T5 Raffel et al. (2020). SLUL: Sign Language Understanding Linker. SLP-MoE: Sign Language Production Mixture-of-Experts. SAGM Loss: Semantic-Aware Gloss Masking Loss. The results show that all the changes were beneficial.
@@ 3.1 Experimental Setup
As mentioned above, our training is based on the Prompt2Sign dataset Fang et al. (2025a) and the American Sign Language (ASL) portion of WLASL WLASL. They have 30k and 10k video clips of ASL respectively, and some of the missing parts (Prompts and Glosses) were properly completed or deleted.
@@ 3.2 Evaluation for Sign Language Production
In Table 3, we present final video generation results with SLP-MoE. Compared to Fast-SLP Transformers (SignDiffusion Fang et al. (2025c)), we achieve 48.6% improvement on BLEU-4 (dev) and 63.7% on test set. Our test BLEU-4 (21.03) surpasses latest Neural Sign Actors (13.12) Baltatzis et al. (2024a) by a large margin, validating our hierarchical generation approach.
@@
Neural Sign Actors Baltatzis et al. (2024a) 13.12 47.55
Fast-SLP Transformers Fang et al. (2025c) 17.19 55.85 12.85 46.89
@@
Ablation Study. We conduct ablation studies on the How2Sign dataset to validate each component’s contribution, as shown in Table 2. We evaluate progressive improvements from base T5 Raffel et al. (2020) through SLUL, SAGM Loss, and SLP-MoE additions on both development and test sets.
@@
Zhang Zhang et al. (2025a) - - 27.60 66.40
SignLLM Fang et al. (2025a) 23.10 58.76 22.05 56.46
@@
We compare SLUL’s Prompt2Gloss performance with previous Text2Gloss approaches in Table 4. This is important as users naturally input prompts like “How do you sign ‘I carried it’?” rather than simplified text. Our approach achieves test BLEU-4 of 30.74 and ROUGE of 69.72, representing 43.3% and 23.5% improvements over SignLLM Fang et al. (2025a). Despite handling more complex prompts than baseline Text2Gloss tasks, we outperform all methods including Zhang et al. (2025a) by 11.4% on test BLEU-4. This demonstrates that SAGM Loss effectively handles prompt nuances including ambiguity and implicit semantics. Consistent performance across dev and test sets indicates robust generalization rather than overfitting.
@@ 3.3 Evaluation for Video Generation Quality
The simplified pose targets shown in Figure 5(a) demonstrate a critical innovation of our approach—the model automatically learns to generate clean, semantically meaningful pose representations through the synergistic integration of rule-based pose priors and SLUL-based semantic guidance. These simplified poses effectively eliminate the noise, jitter, and temporal instability commonly present in raw pose estimations from conventional video analysis tools such as DWPose Yang et al. (2023). Of course, what is even more important is as shown in the first row of the picture: different sign language users have different expressions. These gestures are not aligned in time and space, which causes the models that previously learned average gestures to often be unstable, or to generate restricted or unconfident gestures. However, our method will automatically learn deterministic gestures. As long as one of the multiple gestures is completed, it is considered correct. Therefore, our generation part is more stable.
@@
Stoll et al. Stoll et al. (2020) 0.727 0.533 23.17 64.01
SignGAN Saunders et al. (2022) 0.759 0.605 22.05 27.75
SignDiff Fang et al. (2025c) 0.849 0.676 20.04 25.22
@@
Zhang Zhang et al. (2025a) 0.864 - - 56.28
SignDiff Fang et al. (2025c) 0.849 0.676 20.04 25.22
@@
Figure 6: Ablation Study on SLP-MoE’s Training Efficiency: We examined the efficiency under the condition where there was no SLP-MoE (using the regular Pose2Vid model instead Cheng et al. (2025)). The results showed that using our SLP-MoE for joint training could achieve more thorough convergence, train faster, and produce better results.
@@ A.1 Sign Language Production
Sign Language Production (SLP) aims to generate sign language videos or motion sequences from spoken language inputs, bridging communication gaps for the deaf and hard-of-hearing community Shuler et al. (2013). Early deep learning approaches decomposed SLP into cascaded steps: Text-to-Gloss (T2G) Abdullah et al. (2025), Gloss-to-Pose (G2P) Xie et al. (2023), and Pose-to-Sign (P2S) Moryossef et al. (2025). However, these pipeline methods struggled with smooth motion blending between signs. Recent works have explored end-to-end approaches using transformer architectures and diffusion models. Neural Sign Actors Baltatzis et al. (2024b) proposed a diffusion-based model trained on 4D signing avatars with anatomically informed graph neural networks defined on the SMPL-X skeleton for realistic 3D sign production. T2S-GPT Yin et al. (2024) introduced dynamic vector quantization for autoregressive SLP, while SignLLM Fang et al. (2025a) extended SLP to a multilingual framework supporting eight sign languages through novel reinforcement learning components. Despite these advances, generating realistic, semantically accurate signing motions with fine-grained hand and facial expressions remains challenging walsh2025slrtp2025signlanguageproduction; TII-Web-2023-Falcon; Cho-arxiv-2022-Prompt; Fang et al. (2025b); Hwang et al. (2021).
@@ A.2 Motion Generation
Human motion generation has witnessed significant progress with the adoption of diffusion models. MotionDiffuse Zhang et al. (2022) was among the first diffusion-based frameworks for text-driven motion generation, demonstrating probabilistic mapping through denoising steps and enabling fine-grained body part control. MDM (Motion Diffusion Model) Tevet et al. (2022) introduced a transformer-based architecture with classifier-free guidance, predicting motion samples rather than noise to facilitate geometric losses on joint positions and velocities. LGTM Sun et al. (2024) proposed a Local-to-Global pipeline leveraging large language models to decompose motion descriptions into part-specific narratives, addressing local semantic accuracy issues. Recent works have also explored music-to-dance synthesis, compositional action generation Liu et al. (2024), and stylized motion control Zhong et al. (2025). These advances provide foundational techniques for generating temporally coherent human movements, which are directly applicable to sign language motion synthesis where precise hand articulation and body coordination are essential.
Figure 7: The framework of Wan-Animate Cheng et al. (2025). It utilizes a unified input paradigm and a decoupled control strategy. Spatially-aligned skeleton signals are directly injected into the DiT backbone for body control, while facial expressions are driven by implicit features extracted from reference images and injected via cross-attention layers.
@@ A.3 Controllable Video Generation
Controllable video generation has emerged as an active research direction, extending image-based control mechanisms to the temporal domain. ControlNet Zhang et al. (2023a) pioneered the integration of structural controls such as depth maps, edge detection, and human pose into text-to-image diffusion models through zero convolution layers. ControlVideo Zhang et al. (2023b) adapted ControlNet to video generation without fine-tuning, inheriting high-quality consistent frame generation through fully cross-frame interaction and interleaved-frame smoothing. ControlNeXt Peng et al. (2025) further improved efficiency by reducing trainable parameters up to 90% while supporting diverse conditional controls including pose sequences for video diffusion. Follow-Your-Pose Ma et al. (2024) and MimicMotion Zhang et al. (2025b) specifically addressed human video generation with pose guidance. Most recently, Wan-Animate Cheng et al. (2025) advances this field by leveraging a Diffusion Transformer (DiT) Peebles and Xie (2023); Tevet et al. (2022); Fang et al. (2025d); Saharia et al. (2022) architecture. As illustrated in Figure 7, it employs a decoupled control paradigm that integrates spatially-aligned skeleton signals for body movements and implicit facial features via cross-attention for fine-grained expression reenactment. These controllable generation techniques offer promising directions for sign language video synthesis SLcomprehensibility; signor; Xu-SIGPLAN-2022-Systematic, where pose-conditioned generation can ensure anatomically correct and semantically aligned signing motions.